[Ncep.nhc.nco_contacts] HSU Shift Log for Thu Jan 14 2016 06:00:38 GMT-0500 (EST)

HSU Shift Log michael.j.brennan at noaa.gov
Thu Jan 14 11:00:44 UTC 2016


HSU Shift Log for Thu Jan 14 2016 06:00:38 GMT-0500 (EST)

Shift: A

Forecaster: SRS

Technical issues: No SHIPS output due to OFCI 72-h point being north of  
60N. TWO graphic continues to show "X" along with TS symbol. ATCF satellite  
fix importer still not working due to Alex still being carried as AL902016  
instead of AL012016; TAFB corrected this.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160114/8566f348/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list